Some systems are experiencing issues.

ETH-Services Status

Status of AS200462

  • Network

    Last updated 2025-02-12 00:43:47

    Operational
  • Power

    Last updated 2025-02-12 00:43:47

    Operational
  • Cooling

    Last updated 2025-02-12 00:43:47

    Operational
  • Cogent / AS174

    Last updated 2025-02-12 00:43:47

    Operational
  • Arelion / AS1299

    Last updated 2025-02-12 00:43:47

    Operational
  • NTT / AS2914

    Last updated 2025-02-27 23:17:11

    Operational
  • Inter.Link / AS5405

    Last updated 2025-04-02 23:16:53

    Performance Issues
  • Orange / AS5511

    Last updated 2025-02-12 00:43:47

    Operational
  • Liberty Global / AS6830

    Last updated 2025-02-12 00:43:47

    Operational
  • RETN / AS9002

    Last updated 2025-02-16 14:09:53

    Operational
  • Meerfarbig / AS34549

    Last updated 2025-03-04 14:17:37

    Operational
  • Cloud

    Last updated 2025-02-12 00:43:47

    Operational
  • DNS

    Last updated 2025-02-12 00:43:47

    Operational
  • DDoS-Filter

    Last updated 2025-02-12 00:43:47

    Operational
  • E-Mail

    Last updated 2025-02-12 00:43:47

    Operational
  • VServer

    Last updated 2025-02-12 00:43:47

    Operational
  • VServer - Backups

    Last updated 2025-04-01 14:22:25

    Operational
  • Web

    Last updated 2025-02-12 00:43:47

    Operational

Past Incidents

Fixed

Fixed

3 weeks ago —

The incident has been resolved.

Watching

3 weeks ago —

Arelion has installed various reroutes. Sessions have been re-enabled and no further impact is expected. We will contiue to monitor within the next hours.

3 weeks ago —

We are currently observing high latency and packetloss via Arelion towards various targets in Italy, mostly Milan. To avoid further impact, sessions have been drained.

According to Arelion NOC, the problem is due to outage in Milan. We are awaiting further clarification.

Fixed

Fixed

3 weeks ago —

Cogent NOC has taken steps to avoid impact on their transport. Path priotistation has been restored.

Identified

3 weeks ago —

The issue remains unresolved. Cogents process of TE is ongoing.

To avoid further delay in reactivation of the upstream, any routes learned from Cogent in NA have been deprioritized. No customer impact is expected.

3 weeks ago —

We are currently observing low throuput and packetloss via Cogent towards US targets. To avoid further impact, sessions have been drained. A ticket with the carrier was openend and we are currently awaiting clarification.